# Ornstein-Uhlenbeck fluctuations for the line counting process of the ancestral selection graph

### v2

@@ 1 Introduction
In particular, in the case of the Moran model with selection the ASG is in pathwise duality with the forward in time frequency process. The event that a randomly chosen individual at time $t$ is of the beneficial type is the same as the event that an individual of the beneficial type is among the potential ancestors. If one supposes that a beneficial mutation arises at time 00 with equal probability in any present individual, then the probability that this mutant is among the potential ancestors is proportional to the number of potential ancestors. Therefore, the fluctuations in the number of potential ancestors determine how the probability that a randomly sampled individual at time $t$ is of the beneficial type differs depending on when the mutation arises.
@@ 2.1 The Moran model and the ancestral selection graph
Throughout our work, we consider the classical two-type Moran model [24] in continuous-time with directional selection, constant population size $N$, which is often envisioned as a graph as depicted in Figure 1, also see [10, Chapter 6]. The $N$ individuals are labelled by $[N]:=\{1,\dots,N\}$, each individual $i\in[N]$ is represented as a line from left (past) to the right (future). For each individual $i\in[N]$ we consider a Poisson Point process $P^{i}$ (governing the neutral reproduction) with rate $\frac{\gamma}{2}>0$ and assume $(P^{i},i\in[N])$ to be independent. At each point $t^{i}$ of $P^{i}$ individual $i$ reproduces. At reproduction individual $i$ samples an uniform individual $j\in[N]$ (possibly $i$ itself) and puts its offspring there and individual $j$ inherits the type of $i$. Consequently, we draw an arrow pointing from $i$ to $j$ at time $t^{i}$ in the graphical construction. This leads to the classical graphical construction of the neutral Moran model.
@@
Figure 1: Graphical representation of the Moran model with selection and $N=5$ individuals. Selective arrows are depicted as dashed arrows. Initially, only individual $3$ is of the beneficial type, shown in blue. Individual $3$ is not replaced at time $t_{0}$ since individual $5$ is of wild-type. At time $t_{1}$ the beneficial descendant (individual $4$) of individual $3$ is able to use the selective arrow for reproduction.
@@
Figure 2: In this picture we are following the potential ancestry of individual $2$ from right to left. At time $t_{1}$ the process encounters a selective event, hence the ancestral graph splits and individual $4$ is considered a potential ancestor. At time $t_{0}$ there is coalescence event happening since individual $4$ chooses individual $3$ as a parent, hence there is one line less to follow. In the end individuals $\{1,3,5\}$ are considered potential ancestors of individual $2$. Hence, $\mathcal{A}^{N}_{0}=\{1,3,5\}$ and $B^{N}_{0}=3$.
@@ 2.2 Logistic Branching Processes
The line counting process $B^{N}$ of the ancestral selection graph belongs to the class of logistic branching processes in continuous-time, if one allows frequency dependence in the branching rate. The logistic branching process as in [21] does not have a frequency-dependent branching rate and instead assumes that the branching events are independent for each individual. Inspired by the definition of a logistic branching process in [21], we define a logistic branching process with frequency-dependent branching and refer to it simply as the logistic branching process.
@@
Thus, $\rho$ represents the total birth rate of an individual, whereas the function $h$ describes the frequency dependence in the branching rate. In addition, each individual dies independently at rate $d\geq 0$. The competition between individuals is modelled as a pairwise interaction between individuals, resulting in an additional death rate per individual of $c(k-1)$, with $c\geq 0$. Let $(X_{t},t\geq 0)$ denote the number of individuals alive at time $t$. The transition matrix $Q=(q(i,j)_{i,j\in\mathbb{N}_{0}})$ of $X$ is given by
$q(i,j)=\begin{cases}\rho ih(i)\pi_{j-i},\quad&\text{if }i\geq 1\text{ and }j>i,\\ di+ci(i-1),\quad&\text{if }i\geq 1\text{ and }j=i-1,\\ -i(d+\rho h(i)+c(i-1)),\quad&\text{if }i\geq 1\text{ and }j=i,\\ 0,\quad&\text{otherwise.}\end{cases}$ (6)
@@
We obtain the line counting process of the Moran model by choosing $\rho=s_{N}$, $\pi_{1}=1$, $h(k)=1-k/N$, $c=\gamma/(2N)$ and $d=0$.
@@ 2.3.1 Convergence for the logistic branching process
where $(W_{t},t\geq 0)$ is a standard Brownian motion independent of $Y$.
@@
$h_{N}(\mu^{X}_{N})-1=\mu^{X}_{N}/N=\frac{2}{\gamma}s_{N}\stackrel{{\scriptstyle!}}{{=}}o(\sigma^{X}_{N}/\mu^{X}_{N})=o(s_{N}^{-\frac{1}{2}}N^{\frac{1}{2}}).$ (14)
@@ 2.3.2 Convergence of the line counting process
The next theorem concerns the limiting distribution of $\widetilde{B}^{N}$, as $N\to\infty$. In the Theorem we are dealing with the regime of strong selection, that is $s_{N}\to s>0$ as $N\to\infty$ and with the regime of moderate selection, which assumes
@@
[…]
where $(W_{t},t\geq 0)$ is a standard Brownian motion independent of $Y$.
@@
  • The case of weak selection, i.e. $s_{N}N\to\alpha\in(0,\infty)$, is excluded by Theorem 2 and in fact it is no longer true. Instead, by the same method as in the present work one can show that
@@ 2.4 Martingale problems and convergence of generators
[…]
@@
It is hardly possible to verify condition (27) for the proof of Theorem 1 or Theorem 2. In particular, the convergence of generators is no longer valid in regions far away from the centre of attraction $\mu_{N}^{X}$ or $\mu_{N}^{B}$ as the drift towards these centers grows faster than linear far away form the center. Therefore, the relaxation of (27) is needed. A similar result can be found in the lecture notes [17] (in German). The result there is more general in the sense that in addition it covers the case of discrete time Markov chains, whereas in our case we are covering the continuous-time case. The proof in [17] is obtained from scratch without relying on the methods of [12], making use of these methods allows us to shorten the proof. However, we want to stress that the knowledge of the results in [17] has helped us tremendously in order to provide a short proof of the statement.
@@ 3 Proof of the main results
[…]
@@
[…]
@@
[…]
@@
Now as $n\to\infty$ the first term converges to 00 by dominated convergence and assumption (25) thus for $n$ large enough the above can be made smaller than any $\varepsilon^{\prime}>0$ by choosing $\varepsilon$ small enough. Consequently, the term in (45) converges to 00 as $n\to\infty$ and the claim follows by an application of Corollary 8.15 in [12]. ∎
@@ 3.2 Proof of Theorem 1
$=-\bar{\pi}x_{N}f_{N}^{\prime}(x_{N})+\mathcal{O}(\sigma_{N}^{-1})+o(1).$
Similarly, we get for the terms involving the second derivative of $f_{N}$
@@
In fact we will prove a slightly stronger statement, that is for some $0\leq\lambda_{1}\leq\lambda_{2}<\infty$ and all $x\in[-\lambda_{1},\lambda_{1}]$ it holds for all $N\geq N_{0}\in\mathbb{N}$
@@
As a first step we estimate the maximal number of transitions $\widetilde{X}^{N}$ can have while staying in $[-\lambda_{2},\lambda_{2}]$ for a time interval of length $T$. For that reason let $E^{N}_{1},E^{N}_{2},\dots$ be i.i.d. standard exponential random variables with rate
@@
where $q(x,x)$ is as in (6). Hence, $r_{N}$ is the maximal jump rate of $\widetilde{X}^{N}$ in $[-\lambda_{2},\lambda_{2}]$.
For convenience in notation, let $x^{+}=\sup\{[\mu_{N}-\lambda_{2}\sigma_{N},\mu_{N}+\lambda_{2}\sigma_{N}]\cap\mathbb{N}\}$, then
@@
for some finite constant $C>0$. Therefore, the probability that $\widetilde{X}^{N}$ has more jumps than $m_{N}$ before time $T$, if
@@
Note that the function $y\mapsto y-1-\ln(y)>0$ for all $y\in(0,1)$ and by assumption (10) we get $m_{N}\to\infty$. Hence, the probability in (57) tends to 00 as $N\to\infty$.
@@
$\leq\frac{4v^{2}(1+\delta)TC}{(\lambda_{2}-\lambda_{1})^{2}}+\mathbb{P}(A_{m_{N}}^{c}),$ (60)
@@ 3.3 Auxiliary Lemma
In $K_{\eta}$ it is difficult to estimate the behaviour of $M^{N}$ (as seen in Lemma 3.4), which is the reason why we work in the bigger boxes $K_{c}$ for $c>\eta$. As a next step we need to consider the initial value $x\in K_{\eta}$.
If $x\in K_{\eta}$, we wait until $M^{N}$ leaves $K_{\eta}$, where on the event $H_{n}=\{|M_{k}^{N}-M_{k-1}^{N}|\leq\eta\sigma_{N},\forall\,k\leq n\}$ $M^{N}$ is in $K_{2\eta}$ after leaving $K_{\eta}$. In particular, we can conclude that for $x\in K_{\eta}$ we have with $x^{-}=\lfloor\mu_{N}-2\eta\sigma_{N}\rfloor$, $x^{+}=\lceil\mu_{N}+2\eta\sigma_{N}\rceil$
@@
where $p_{x}$ denotes the probability that $M^{N}$ leaves $K_{\eta}$ upwards. Hence, we consider the case $x\in K_{2\eta}$ for estimating the probability in (68) and consider the other case later.
By the construction of $K_{2\eta}$, the process $M^{N}$ has a drift towards $\mu_{N}$, while being in $K_{2\eta}\setminus K_{\eta}$. Hence, the probability that $M^{N}$ leaves the box $[\mu_{N}\pm\eta^{\prime}\sigma_{N}]\cap\mathbb{N}$ can be estimated from above by the probability that an unbiased random walk $(S^{N}_{k},k\geq 0)$ leaves the box $[\mu_{N}\pm(\eta^{\prime}-2\eta)\sigma_{N}]$, where $S^{N}$ is specified in the next paragraph. $S^{N}$ is chosen such that whenever $M_{N}$ is in $K_{j\eta}\setminus K_{2\eta}$ for $j>2$ the probability to deviate from $\mu_{N}$ by more than $\eta^{\prime}\sigma_{N}$ is higher for $S_{N}$ and in addition whenever $M^{N}$ would enter $K_{2\eta}$ the process $S^{N}$ is set to some state $x\notin K_{2\eta}$, hence $S^{N}$ stays closer to the boundary of $K_{\eta^{\prime}}$.
Let $S^{N}=(S^{N}_{n},n\geq 0)$ be a random walk with jump distribution $\tilde{p}^{N}(x,y)$, which is obtained from $p^{N}(x,y)$ by alternating the probability to jump downwards such that $\tilde{p}^{N}(x,y)$ is mean zero for all $x\in\mathbb{N}$. By this construction we ensure that $S^{N}$ can be coupled to $M^{N}$ such that whenever both of them are in the same state space $x\notin K_{2\eta}$, $M^{N}$ jumps with higher probability in the direction of $\mu_{N}$.
In addition we “cut out” $K_{2\eta}$ from the state space of $S^{N}$ by saying that whenever $M^{N}$ would jump from above into $K_{2\eta}$ we put $S^{N}$ to the lower boundary of $K_{2\eta}$ (meaning the natural number smaller than $\inf K_{2\eta}$), hence $S^{N}$ jumps over $K_{2\eta}$. In addition, whenever $M^{N}$ would jump into $K_{2\eta}$ from below, we know that on $H_{n}$, $M^{N}$ can’t jump higher than $K_{\eta}$. Say $S^{N}$ would jump to a height $d$ above $\inf K_{2\eta}$ we put $S^{N}$ at height $d+\sup K_{2\eta}$. This construction ensures that $S^{N}$ is an unbiased random walk on $\mathbb{Z}\cap K_{2\eta}^{c}$. See also Figure 3 for an illustration.
@@
The variance estimate is simple now by simply considering the maximal variance of $\tilde{p}^{N}(x,\cdot)$ for $x\in K_{\eta^{\prime}}$, we have
@@
Let us consider the remaining case, where $x\notin K_{2\eta}$ but $x\in K_{\lambda}$ as assumed in the Lemma. Here the argument works analogously, we estimate
@@
Lastly, we have $x\in K_{2\eta},x\notin K_{\eta}$. Here we know that there exists a drift towards $\mu_{N}$ but we cannot guarantee that the process does not jump immediately out of $K_{2\eta}$. However, this is no problem at all, since we did not condition $S^{N}$ to have these small jumps. In particular the probability that $M^{N}$ escapes $K_{\eta^{\prime}}$ is smaller than the probability that $S^{N}$ does in time $n$, hence we obtain the same bound as (76) with $\lambda=2\eta$.
@@
$\mathbb{P}(H_{n}^{c})\leq\mathbb{P}(\cup_{0\leq k\leq n}\{X_{j}>\sigma_{N}\})\leq n\mathbb{P}(X_{1}>\sigma_{N})\leq\frac{nC}{\sigma_{N}^{3}},$ (78)

### v3

@@ 1 Introduction
In particular, in the case of the Moran model with selection the ASG is in pathwise duality with the forward in time frequency process. The event that a randomly chosen individual at time $t$ is of the beneficial type is the same as the event that an individual of the beneficial type is among the potential ancestors. If one supposes that a beneficial mutation arises at time 0 with equal probability in any present individual, then the probability that this mutant is among the potential ancestors is proportional to the number of potential ancestors. Therefore, the fluctuations in the number of potential ancestors determine how the probability that a randomly sampled individual at time $t$ is of the beneficial type differs depending on when the mutation arises.
@@ 2.1 The Moran model and the ancestral selection graph
Throughout our work, we consider the classical two-type Moran model [24] in continuous-time with directional selection, constant population size $N$, which is often envisioned as a graph as depicted in Figure 1, also see [10, Chapter 6]. The $N$ individuals are labelled by $N]:=\{1,\dots,N\}$, each individual $i\in[N]$ is represented as a line from left (past) to the right (future). For each individual $i\in[N]$ we consider a Poisson Point process $P^{i}$ (governing the neutral reproduction) with rate $\frac{\gamma}{2}>0$ and assume $(P^{i},i\in[N])$ to be independent. At each point $t^{i}$ of $P^{i}$ individual $i$ reproduces. At reproduction individual $i$ samples an uniform individual $j\in[N]$ (possibly $i$ itself) and puts its offspring there and individual $j$ inherits the type of $i$. Consequently, we draw an arrow pointing from $i$ to $j$ at time $t^{i}$ in the graphical construction. This leads to the classical graphical construction of the neutral Moran model.
@@
Figure 1: Graphical representation of the Moran model with selection and $N=5$ individuals. Selective arrows are depicted as dashed arrows. Initially, only individual $3$ is of the beneficial type, shown in blue. Individual 33 is not replaced at time $t_{0}$ since individual $5$ is of wild-type. At time $t_{1}$ the beneficial descendant (individual 44) of individual $3$ is able to use the selective arrow for reproduction.
@@
Figure 2: In this picture we are following the potential ancestry of individual $2$ from right to left. At time $t_{1}$ the process encounters a selective event, hence the ancestral graph splits and individual $4$ is considered a potential ancestor. At time $t_{0}$ there is coalescence event happening since individual 44 chooses individual 33 as a parent, hence there is one line less to follow. In the end individuals $\{1,3,5\}$ are considered potential ancestors of individual 22. Hence, $\mathcal{A}^{N}_{0}=\{1,3,5\}$ and $B^{N}_{0}=3$.
@@ 2.2 Logistic Branching Processes
The line counting process $B^{N}$ of the ancestral selection graph belongs to the class of logistic branching processes in continuous-time with a frequency-dependent branching rate. The logistic branching process as it is commonly defined in [21] does not have a frequency-dependent branching rate and instead assumes that the branching events are independent for each individual. Inspired by this definition of a logistic branching process in [21], we allow a logistic branching process to have a frequency-dependent branching rate, but refer to it simply as the logistic branching process.
@@
Thus, $\rho$ represents the total birth rate of an individual, whereas the function $h$ describes the frequency dependence in the branching rate. In addition, each individual dies independently at rate $d\geq 0$. The competition between individuals is modelled as a pairwise interaction between individuals, resulting in an additional death rate per individual of c(k1)c(k-1), with $c\geq 0$. Let $(X_{t},t\geq 0)$ denote the number of individuals alive at time $t$. The transition matrix $Q=(q(i,j)_{i,j\in\mathbb{N}_{0}})$ of $X$ is given by
[…]
@@
We obtain the line counting process of the Moran model by choosing $\rho=s_{N}$, $\pi_{1}=1$, h(k)=1k/Nh(k)=1-k/N, $c=\gamma/(2N)$ and $d=0$.
@@ 2.3.1 Convergence for the logistic branching process
where $(W_{t},t\geq 0)$ is a standard Brownian motion independent of $Y_{0}$.
@@
$h_{N}(\mu^{X}_{N})-1=\mu^{X}_{N}/N=\frac{2}{\gamma}s_{N}$ (14)
to be at most
$o(\sigma^{X}_{N}/\mu^{X}_{N})=o(s_{N}^{-\frac{1}{2}}N^{-\frac{1}{2}}).$ (15)
@@ 2.3.2 Convergence of the line counting process
The next theorem concerns the functional limit theorem for $B^{N}$, as $N\to\infty$. In the Theorem we are dealing with the regime of strong selection, that is $s_{N}\to s>0$ as $N\to\infty$ and with the regime of moderate selection, which assumes
@@
$dY_{t}=-Y_{t}dt+\sqrt{2}dW_{t},\quad Y_{0}\overset{d}{=}\nu,$ (20)
where $(W_{t},t\geq 0)$ is a standard Brownian motion independent of $Y_{0}$.
@@
  • The case of weak selection, i.e. $s_{N}N\to\alpha\in(0,\infty)$, is excluded by Theorem 2 and in fact the conclusions of Theorem 2 are no longer true. Instead, by the same method as in the present work one can show that
@@ 2.4 Martingale problems and convergence of generators
$M_{f}(t)=f(Y_{t})-\int_{0}^{t}Gf(Y_{s})ds,\qquad Y_{0}\overset{d}{=}\nu,$ (23)
is a $\mathbb{P}$-martingale with respect to the natural filtration $\mathcal{F}_{t}=\sigma(Y_{s},s\leq t)$. A martingale problem $(G,\mathcal{D})$ is unique if for two solutions $(Y,\mathbb{P})$ and $(X,\mathbb{Q})$ with $Y_{0}\overset{d}{=}X_{0}\overset{d}{=}\nu$ it holds
@@
It is hardly possible to verify condition (28) for the proof of Theorem 1 or Theorem 2. In particular, the convergence of generators is no longer valid in regions far away from the centre of attraction $\mu_{N}^{X}$ or $\mu_{N}^{B}$ as the drift towards these centers grows faster than linear far away from the center. Therefore, the relaxation of (28) is needed. A similar result can be found in the lecture notes [17] (in German). The result there is more general in the sense that in addition it covers the case of discrete time Markov chains, whereas in our case we are covering the continuous-time case. The proof in [17] is obtained from scratch without relying on the methods of [12], making use of these methods allows us to shorten the proof. However, we want to stress that the knowledge of the results in [17] has helped us tremendously in order to provide a short proof of the statement.
@@ 3 Proof of the main results
$\lim_{n\to\infty}\mathbb{E}\left[\sup_{t\in\mathbb{Q}\cap[0,T]}|\xi^{n}_{t}-f(Y^{n}_{t})|\right]=0,\quad\sup_{n\geq 1}\mathbb{E}\left[\left(\int_{0}^{T}|\varphi^{n}_{s}|^{p}ds\right)^{\frac{1}{p}}\right]<\infty,$ (31)
@@
$\mathbb{P}\left(\tau_{n}<T\right)$ $=\mathbb{P}\left(\int_{0}^{t}|g_{n}(Y^{n}_{s})|^{2}ds\geq t(\lVert g\rVert_{\infty}^{2}+1),\text{ for some }t\in[0,T]\right)$ (32)
@@
$\sup_{n\geq 1}\mathbb{E}\left[\left(\int_{0}^{T}|\varphi^{n}_{s}|^{2}dx\right)^{\frac{1}{2}}\right]\leq T^{\frac{1}{2}}(\lVert g\rVert_{\infty}^{2}+1)^{\frac{1}{2}}<\infty.$ (39)
@@
Now as $n\to\infty$ the first term converges to 0 by dominated convergence and assumption (26) thus for $n$ large enough the above can be made smaller than any $\varepsilon^{\prime}>0$ by choosing $\varepsilon$ small enough. Consequently, the term in (46) converges to 0 as $n\to\infty$ and the claim follows by an application of Corollary 8.15 in [12]. ∎
@@ 3.2 Proof of Theorem 1
$={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tfrac{1}{\sigma_{N}}f_{N}^{\prime}(x_{N})\bigg[(\mu_{N}+x_{N}\sigma_{N})\bar{\pi}+o(\sigma_{N})-o(c_{N}\mu_{N}\rho_{N}^{-1})-}$
$\qquad\qquad{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}-\frac{c_{N}}{\rho_{N}}\mu_{N}^{2}-2\frac{c_{N}\mu_{N}x_{N}\sigma_{N}}{\rho_{N}}+\mathcal{O}(c_{N}\sigma_{N}^{2}\rho_{N}^{-1})+\mathcal{O}(c_{N}\mu_{N}))\bigg]}$
@@
$=-\bar{\pi}x_{N}f_{N}^{\prime}(x_{N})+\mathcal{O}(\sigma_{N}^{-1})+o(1),$
where for the first equality we have used (11) and that $d_{N}=o(c_{N})$, see (10). Similarly, we get for the terms involving the second derivative of $f_{N}$
@@
In fact we will prove a slightly stronger statement, that is for some $0\leq\lambda_{1}\leq\lambda_{2}<\infty$, such that for all $x\in[-\lambda_{1},\lambda_{1}]$ there exists some $N_{0}\in\mathbb{N}$ such that for $N\geq N_{0}\in\mathbb{N}$ it holds that
@@
As a first step we estimate the maximal number of transitions $\widetilde{X}^{N}$ can have while staying in $\lambda_{2},\lambda_{2}]$ for a time interval of length $T$. For that reason let $E^{N}_{1},E^{N}_{2},\dots$ be i.i.d. standard exponential random variables with rate
@@
where $q(x,x)$ is as in (6). Hence, $r_{N}$ is the maximal jump rate of $\widetilde{X}^{N}$ in $\lambda_{2},\lambda_{2}]$.
For convenience in notation, let $x^{+}=\sup\{[\mu_{N}-\lambda_{2}\sigma_{N},\mu_{N}+\lambda_{2}\sigma_{N}]\cap\mathbb{N}\}$, then by bounding the transition rates in (6)
@@
for some finite constant $C>0$, where we also used that $c_{N}(x^{+})^{2}/\rho_{N}=\mathcal{O}(c_{N}\mu_{N}^{2}/\rho_{N})=\mathcal{O}(\mu_{N})$, which follows from (9) and (10). Therefore, $\widetilde{X}^{N}$ has more jumps than $m_{N}$ before time $T$, if
@@
Note that the function $y\mapsto y-1-\ln(y)>0$ for all $y\in(0,1)$ and by assumption (10) we get $m_{N}\to\infty$. Hence, the probability in (58) tends to 0 as $N\to\infty$.
@@
$\leq\frac{2v^{2}(1+\delta)TC+2v^{2}}{(\lambda_{2}-\lambda_{1})^{2}}+\mathbb{P}(A_{m_{N}}^{c}),$ (61)
@@ 3.3 Auxiliary Lemma
In $K_{\eta}$ it is difficult to estimate the behaviour of $M^{N}$ (as seen in Lemma 3.4), which is the reason why we work in the bigger boxes $K_{c}$ for $c>\eta$. As a next step we do a case distinction for the initial value $x\in K_{\lambda}$ and distinguish the cases i) $x\in K_{\eta}$, ii) $x\in K_{2\eta}\setminus K_{\eta}$, iii) $x\in K_{\lambda}\setminus K_{2\eta}$.
Case i) If $x\in K_{\eta}$, we wait until $M^{N}$ leaves $K_{\eta}$, where on the event $H_{n}=\{|M_{k}^{N}-M_{k-1}^{N}|\leq\eta\sigma_{N},\forall\,k\leq n\}$ $M^{N}$ is in $K_{2\eta}$ after leaving $K_{\eta}$. In particular, we can conclude that for $x\in K_{\eta}$ we have with $x^{-}=\lfloor\mu_{N}-2\eta\sigma_{N}\rfloor$, $x^{+}=\lceil\mu_{N}+2\eta\sigma_{N}\rceil$
@@
where $p_{x}$ denotes the probability that $M^{N}$ leaves $K_{\eta}$ upwards. Hence, we consider the case $x\in K_{2\eta}$ for estimating the probability in (69).
Case ii) If $x\in K_{2\eta}\setminus K_{\eta}$, by the construction of $K_{2\eta}$, the process $M^{N}$ has a drift towards $\mu_{N}$, while being in $K_{2\eta}\setminus K_{\eta}$. Hence, the probability that $M^{N}$ leaves the box $\mu_{N}\pm\eta^{\prime}\sigma_{N}]\cap\mathbb{N}$ can be estimated from above by the probability that an unbiased random walk $(S^{N}_{k},k\geq 0)$ leaves the box $\mu_{N}\pm(\eta^{\prime}-2\eta)\sigma_{N}]$, where $S^{N}$ is specified in the next paragraph. $S^{N}$ is chosen such that whenever $M_{N}$ is in $K_{j\eta}\setminus K_{2\eta}$ for $j>2$ the probability to deviate from $\mu_{N}$ by more than $\eta^{\prime}\sigma_{N}$ is higher for $S_{N}$ and in addition whenever $M^{N}$ would enter $K_{2\eta}$ the process $S^{N}$ is set to some state $x\notin K_{2\eta}$, hence $S^{N}$ stays closer to the boundary of $K_{\eta^{\prime}}$.
Let $S^{N}=(S^{N}_{n},n\geq 0)$ be a random walk with jump distribution $\tilde{p}^{N}(x,y)$, which is obtained from $p^{N}(x,y)$ by alternating the probability to jump downwards such that $\tilde{p}^{N}(x,\cdot)$ is mean zero for all $x\in\mathbb{N}$. Note that $S^{N}$ jumps down at most $1$, hence we can simply increase or decrease this probability as necessary. By this construction we ensure that $S^{N}$ can be coupled to $M^{N}$ such that whenever both of them are in the same state space $x\notin K_{2\eta}$, $M^{N}$ jumps with higher probability in the direction of $\mu_{N}$.
In addition we “cut out” $K_{2\eta}$ from the state space of $S^{N}$ by saying that whenever $S^{N}$ would jump from above into $K_{2\eta}$ we put $S^{N}$ to the lower boundary of $K_{2\eta}$ (meaning the natural number smaller than $\inf K_{2\eta}$), hence $S^{N}$ jumps over $K_{2\eta}$. In addition, whenever $S^{N}$ would jump into $K_{2\eta}$ from below, we know that on $H_{n}$, $M^{N}$ can’t jump higher than $K_{\eta}$. Say $S^{N}$ would jump to a height $d$ above $\inf K_{2\eta}$ we put $S^{N}$ at height $d+\sup K_{2\eta}$. This construction ensures that $S^{N}$ is an unbiased random walk on $\mathbb{Z}\cap K_{2\eta}^{c}$. See also Figure 3 for an illustration.
@@
We estimate the variance appearing above by considering the maximal variance of the distribution $\tilde{p}^{N}(x,\cdot)$ for all $x\in K_{\eta^{\prime}}$. We have
@@
Lastly, to finish this case we have to consider all starting values $x\in K_{2\eta},x\notin K_{\eta}$. Here we know that there exists a drift towards $\mu_{N}$ but we cannot guarantee that the process does not jump immediately out of $K_{2\eta}$. However, this is no problem at all, since we did not condition $S^{N}$ to have these small jumps. In particular the probability that $M^{N}$ escapes $K_{\eta^{\prime}}$ is smaller than the probability that $S^{N}$ does in time $n$, hence we obtain the same bound as (76) with $x$ instead of $x^{+}$.
Case iii) Let us consider the remaining case, where $x\in K_{\lambda}$ but $x\notin K_{2\eta}$, as assumed in the Lemma. Here the argument works analogously, we estimate
@@
$\mathbb{P}(H_{n}^{c})\leq\mathbb{P}(\cup_{0\leq k\leq n}\{X_{j}>\eta\sigma_{N}\})\leq n\mathbb{P}(X_{1}>\eta\sigma_{N})\leq\frac{nC}{\eta^{3}\sigma_{N}^{3}},$ (79)